The cost, in millions of dollars, to remove x% of pollutin in a lake modeled by
c = 6,000/200 - 2x.
3rd question: What is the cost to remove 99% of the pollutant?

Answer by stanbon(75887)

The cost, in millions of dollars, to remove x% of pollutin in a lake modeled by
c = 6,000/200 - 2x.
3rd question: What is the cost to remove 99% of the pollutant?
------------------------------
99 = 6000/(200-2x)
99(200-2x) = 6000
200-2x = 60.6060
2x = 139.3930
x = $69.696969... million
x = $69,696,969.69
Cheers,
Stan H.